OpenStudy (anonymous):

The display on a calculator screen is shown below. What number does the calculator display represent? 0.0000000025 2,548,640,000 0.0000000003 254,863,999,999,999.97

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

will fan and medal

OpenStudy (anonymous):

Which shows the expression below simplified? (7 × 10-6) × 0.009 1.6 × 1018 6.3 × 1018 6.3 × 10-8 6.3 × 10-9

OpenStudy (roadjester):

Could you elaborate on what it is you're looking for in this problem?

OpenStudy (anonymous):

just on how to solve it

OpenStudy (roadjester):

But what is the problem we're supposed to solve?

OpenStudy (anonymous):

2nd one

OpenStudy (mathmale):

Double check the first part: (7 x 10 - 6). Did you copy it exactly as it is in your homework assignment? Even if the calculator displays that " x " symbol, " x " does not represent multiplication on the calculator; instead, it represents a variable by the name of "x". Can you take and share a digital image of your homework assignment problems?

OpenStudy (anonymous):

how do i do that

OpenStudy (roadjester):

You can take a picture and upload it.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im on study island

OpenStudy (roadjester):

\[(7\times10^{-6}) \times {0.009}\] I'm assuming this is the problem you need help with, except what are the numbers at the top of the page?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

1.6 × 1018 6.3 × 1018 6.3 × 10-8 6.3 × 10-9

OpenStudy (roadjester):

Then none of the answers provided is correct

OpenStudy (mathmale):

sorry, but I'm not in a position (at the moment) to explain how to take a "screen shot." Do you have a cell phone that could take a photo of your homework, save it to your comoputer, from which you could upload it to openStudy? You could, of course, move on to a new problem. If others are having trouble figuring out what it is that you are asking, perhaps you should start over. Post no more than one problem at a time, include the directions.

OpenStudy (anonymous):

hi mathmale

OpenStudy (anonymous):

sorry I dont

OpenStudy (mathmale):

I doubt that "none of the answers are correct." Please consider following at least one of my suggestions. If at all possible, hold the problems at hand and post a new one, but only ONE at a time.

OpenStudy (anonymous):

ok

OpenStudy (roadjester):

\[0.009\] is equivalent to \[9 \times 10^{-3}\] Therefore you would have \[{7*9} \times {10^{-6}*10^{-3}}\]

OpenStudy (mathmale):

@LGumboymath234: Please let us know if you plan to leave or abandon an OpenStudy conversation such as this one. I note that this conversation has already been closed (though not by me). Thank you.